INFRARED SPECTROSCOPY OF M^{+}(CH_{4})_{n}(H_{2}O)_{3−4} CLUSTERS (M=Li, Na): INDUCING H_{2}O⋯H_{2}O AND H_{2}O⋯CH_{4} HYDROGEN BONDS IN METHANATED CLUSTERS

\underline{OSCAR RODRIGUEZ JR.}, and JAMES M. LISY, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

Experiments on M^{+}(CH_{4})_{n}(H_{2}O)_{3−4} (M=Li and Na) have been carried out using tandem mass spectrometry and infrared spectroscopy in the O-H stretching region from 3400 cm^{−1} to 3800 cm^{−1}. We have found, for example, that a single CH_{4} can induce hydrogen bonded conformers in the cases of Li^{+}(CH_{4})_{1}(H_{2}O)_{3}Ar_{0−1}, which are absent in the Li^{+}(H_{2}O)_{3}Ar_{0−1} spectra^{a}^{,}^{b}. Furthermore, upon addition of multiple CH_{4} ligands, hydrogen bonding is not only maintained, but features associated with H_{2}O⋯CH_{4} hydrogen bonds are more intense, indicating the affinity of CH_{4} to bind to available O-H sites. Spectra of Li^{+}(CH_{4})_{1−3}(H_{2}O)_{4} and Na^{+}(CH_{4})_{1−3}(H_{2}O)_{3−4} clusters will also be discussed, all of which exhibit the curious trait of H_{2}O⋯H_{2}O and H_{2}O⋯CH_{4} hydrogen bonding in the presence of hydrophobic, non-polar CH_{4}. To better understand the nature and onset of H_{2}O⋯CH_{4} hydrogen bonding, the Li^{+}(CH_{4})_{n}(H_{2}O)_{1−2} spectra will be discussed.

aD.J. Miller and J.M. Lisy J. Am. Chem. Soc. 130, 15381 (2008).

bD.J. Miller and J.M. Lisy J. Am. Chem. Soc. 130, 15393 (2008).